# Reversibility in finite-dimensional collapse dynamics

### v1

@@ Abstract
We study finite-dimensional quantum systems with arbitrary, discontinuous collapse events. Formally, we fix a realization map (a physically admissible selector of the collapse dynamics) and do not rely on any regularity of the induced dynamics. We prove the existence of a topologically closed, invariant subset of the projective state space in which any two states can be connected with arbitrarily fine Fubini-Study precision and arbitrarily small integrated energetic cost. This shows that the preservation of information along a realized branch guarantees islands of quasi-reversibility, while genuine irreversibility requires additional ingredients such as non-compactness, explicit erasure, or coupling to reservoirs.
@@ 1.2 The result
The choice of a selector for the dynamics can be interpreted as realizing the extreme, no-information-erasure limit of Landauer’s principle, in the sense highlighted by C.H. Bennett’s resolution of the Maxwell demon paradox ([bennett1982thermodynamics, bennett2003notes, hanson2018landauer]). In particular, we share the following position, expressed by Bennett (from [bennett1982thermodynamics]):
@@
We say that the dynamics $(T_{U,D},\mathcal{E}_{D})$ exhibits an operational arrow of time if there exist states $u],[v]\in\mathbb{P}(\mathcal{H})$ such that
@@
i.e. the energetic effort required by the microscopic law to carry the system from $[u]$ to $[v]$ in forward time is strictly smaller than the minimal energetic effort required to realize the reverse change from $[v]$ back to $[u]$.
@@
We can summarize our main result as follows:
Main result.
In the no–erasure regime, compactness enforces quasi–reversibility: there exists a recurrent, invariant subsystem $S$ allowing quasi-returns with arbitrarily small error and arbitrarily small energetic cost.
In particular, collapse alone does not generate an operational arrow of time on $S$.
Genuine irreversibility, thus, requires additional ingredients: information loss/coarse–graining, non-compact limits, or coupling to reservoirs.
@@ 1.3 Where a naive diagonal grid argument fails
To make this remark more precise, fix a mesh $\epsilon_{k}=2^{-k}$ and choose refining covers $\mathcal{U}_{k+1}\prec\mathcal{U}_{k}$ by $d_{FS}$–balls of radius $\epsilon_{k}$. At scale $k$, scan a long forward segment and pick the first cell $C_{k}\in\mathcal{U}_{k}$ that is visited twice; let $\Gamma_{k}$ be the loop between those visits. This yields strong $\epsilon_{k}$–loops, but the selected cells need not nest: typically $C_{k+1}\nsubseteq C_{k}$. Even after extracting a nested subsequence $C_{k_{r}}\searrow\{x_{\ast}\}$ (via a refinement tree), the discontinuity of $T$ decouples the local dynamics at $x_{\ast}$ from the images traced by the $\Gamma_{k_{r}}$, so neither $x_{\ast}\,\mathcal{SC}_{d_{FS}}\,x_{\ast}$ (where $\mathcal{SC}_{d_{FS}}$ is the strong chain-relation defined in Def. 2.3) nor an invariant, internally $\mathcal{SC}_{d_{FS}}$–transitive set is guaranteed. Summarizing, the grid-refining technique only shows that:
@@ 2.1 Notation and definitions for Quantum Dynamics formalism
where we use the normal units: $\hbar=1.$ If $H$ depends on time, i.e. $H=H(t)$, the evolution is given by the time-ordered exponential
@@ 2.2 Notation and definitions for abstract Topological Dynamics formalism
We say that $(X,f)$ is a topological dynamical system if $X$ is a topological space with $d:X^{2}\to\mathbb{R}_{0}^{+}$ a metric compatible with the topology on $X$ and $f:X\to X$ a map. By $f^{n}(x)$ with $x\in X$ and $n\in\mathbb{N}$ we mean the $n$-times composition $f\circ f\circ\ldots\circ f(x)$. When we say that $(X,f)$ is a compact dynamical system we require, in addition, that $X$ is a compact metric space.
@@ 3 Discrete-time Quantum Dynamics
Note that the existence of a realization map does not conflict with the Kochen–Specker (KS) no-go theorem ([kochen2011problem]). In fact, it addresses a completely different point. KS forbids noncontextual, dispersion-free value assignments to all projections simultaneously (in dimension $\geq 3$), preserving functional relations across incompatible observables. By contrast, throughout we fix a single measurement context $\{P_{j}\}_{j\in\Lambda}$ (one PVM), and the realization map $D:\mathbb{P}(\mathcal{H})\to\Omega$ assigns an outcome sequence only for this context. The assignment is explicitly contextual/history-dependent and makes no claims about outcomes for other, incompatible PVMs. Hence $D$ is not a global valuation on the projection lattice, and the hypotheses of KS are not met.
@@ 4 Reversibility
5 Conclusion
@@
Theorems 4.5 and 4.6 tell us the following fact: in every finite-dimensional, open quantum system, there is a topologically closed, invariant subsystem $S$ where any two states are connected via pseudo-orbits with arbitrarily small Fubini-Study error and energetic cost. By Def. 1.1, this means that $S$ is operationally reversible: no operational arrow of time arises.
@@
It is perhaps useful to frame this result by analogy with the logical structure of classical thermodynamics. The second law provides a universal lower bound on irreversibility, while Carnot’s theorem is an existence result showing that this bound can in fact be saturated: it identifies ideal cycles that achieve reversibility under suitable structural constraints. Landauer’s principle plays a similar role to the second law in the informational setting, furnishing a universal lower bound on dissipation for logically irreversible operations. Our result stands, structurally, in the same relation to Landauer’s principle as Carnot’s theorem does to the second law: it is an existence statement identifying conditions (finite dimensionality and compactness) under which the bound is asymptotically saturated. In this sense, operational quasi-reversibility is enforced within invariant subsystems, in the absence of information erasure or non-compact limits.
@@
Let us finally remark that our argument crucially relies on the retention of the infinite record of outcomes provided by the realization map. This infinite history is essential because, to steer the system toward a target state with arbitrarily small FS error $\epsilon>0$, the perturbations must be tailored to the specific sequence of realized collapse outcomes encountered along the chain. Although only finitely many corrections are required to achieve any fixed $\epsilon$-precision, the number and nature of these corrections generally become unbounded as $\epsilon\to 0$, necessitating, in principle, access to the entire future itinerary to prescribe the appropriate steering strategy at every scale. This echoes Bennett’s resolution of Maxwell’s demon paradox, where he showed that a demon’s finite memory forces logically irreversible erasure to accommodate new data, thereby incurring an entropic cost that preserves the second law ([bennett1982thermodynamics]). Our no-erasure framework mirrors the idealized demon scenario, where an infinite memory sidesteps dissipation. Yet, as Bennett argued for any physical demon, real-world implementations face finite-size constraints that inevitably lead to erasure. Analogously, any practical embodiment of our dynamics with a truncated record would introduce operational irreversibility, which is of course consistent with the second law.
@@
is a bijection between $\lambda$ and the set $\{x^{\gamma}\}_{0\leq\gamma<\lambda}$, so we have that $|\{x^{\gamma}\}_{0\leq\gamma<\lambda}|=|\lambda|$. By Hartogs’ Lemma ([hartogs1915problem]), we can take $\lambda$ so large that $|\{x^{\gamma}\}_{0\leq\gamma<\lambda}|>|X|$, which is a contradiction.

### v2

@@ Abstract
We study finite dimensional quantum systems with arbitrary collapse events, establishing, under no-information-erasure conditions, a structural no-go for trajectory-level operational irreversibility. More precisely, we fix a realization map (a physically admissible selector of the collapse dynamics) and do not rely on any regularity of the induced dynamics. We prove that, for every realization of the collapse dynamics, there exists a topologically closed, forward-invariant subset of the projective state space on which any two states can be connected with arbitrarily fine Fubini–Study precision and arbitrarily small integrated energetic cost. This shows that the preservation of information along a realized branch guarantees islands of quasi-reversibility, while genuine irreversibility requires additional ingredients such as non-compactness, explicit erasure, or coupling to reservoirs.
@@ 1.2 The result
The choice of a selector for the dynamics can be interpreted as realizing the extreme, no-information-erasure limit of Landauer’s principle, in the sense highlighted by C.H. Bennett’s resolution of the Maxwell demon paradox [bennett1982thermodynamics, bennett2003notes, hanson2018landauer]. In particular, we share the following position, expressed by Bennett (from [bennett1982thermodynamics]):
@@
We say that the dynamical law $T_{U,D}$ exhibits an operational arrow of time if there exist states $u],[v]\in\mathbb{P}(\mathcal{H})$ such that
@@
i.e. the minimal energetic effort required to perturb the ambient unitary dynamics so as to carry the system from $[u]$ to $[v]$ in forward time is strictly smaller than the minimal energetic effort required to realize the reverse change from $[v]$ back to $[u]$.
@@
We can summarize our main result as follows.
Main result In the no–erasure regime, compactness enforces quasi–reversibility: given any compact quantum system and any realization of the collapse dynamics, there exists a topologically closed and invariant (with respect to the realization-induced evolution map) subsystem $S$ that is operationally reversible, meaning that, in $S$, quasi-returns with arbitrarily small Fubini-Study error and arbitrarily small energetic cost are allowed between any pair of states.
Genuine irreversibility, thus, requires additional ingredients: information loss/coarse–graining, non-compact limits, or coupling to reservoirs. From a conceptual perspective, this establishes a structural no-go: under finite-dimensional, no-erasure conditions, collapse alone cannot generate an operational arrow of time along individual realized branches.
@@ 1.3 Where a naive diagonal grid argument fails
To make this remark more precise, fix a mesh $\epsilon_{k}=2^{-k}$ and choose refining covers $\mathcal{U}_{k+1}\prec\mathcal{U}_{k}$ by $d_{FS}$–balls of radius $\epsilon_{k}$. At scale $k$, scan a long forward segment and pick the first cell $C_{k}\in\mathcal{U}_{k}$ that is visited twice; let $\Gamma_{k}$ be the loop between those visits. This yields strong $\epsilon_{k}$–loops, but the selected cells need not nest: typically $C_{k+1}\nsubseteq C_{k}$. Even after extracting a nested subsequence $C_{k_{r}}\searrow\{x_{\ast}\}$, the discontinuity of $T$ decouples the local dynamics at $x_{\ast}$ from the images traced by the $\Gamma_{k_{r}}$, so neither $x_{\ast}\,\mathcal{SC}_{d_{FS}}\,x_{\ast}$ (where $\mathcal{SC}_{d_{FS}}$ is the strong chain-relation defined in Def. 2.3) nor an invariant, internally $\mathcal{SC}_{d_{FS}}$–transitive set is guaranteed. Summarizing, the grid-refining technique only shows that:
@@ 2.1 Notation and definitions for Quantum Dynamics formalism
where we use the normal units: $\hbar=1.$ If $H=H(t)$ depends on time, the evolution is given by the time-ordered exponential
@@ 2.2 Notation and definitions for abstract Topological Dynamics formalism
We say that $(X,f)$ is a topological dynamical system if $X$ is a topological space with a metric $d:X^{2}\to\mathbb{R}_{0}^{+}$ compatible with the topology on $X$ and $f:X\to X$ a map. By $f^{n}(x)$ with $x\in X$ and $n\in\mathbb{N}$ we mean the $n$-times composition $f\circ f\circ\ldots\circ f(x)$. When we say that $(X,f)$ is a compact dynamical systemif we require, in addition, that $X$ is a compact metric space.
@@ 3 Discrete-time Quantum Dynamics
Note that the existence of a realization map does not conflict with the Kochen–Specker (KS) no-go theorem [kochen2011problem]. In fact, it addresses a completely different point. KS forbids noncontextual, dispersion-free value assignments to all projections simultaneously (in dimension $\geq 3$), preserving functional relations across incompatible observables. By contrast, throughout we fix a single measurement context $\{P_{j}\}_{j\in\Lambda}$ (one PVM), and the realization map $D:\mathbb{P}(\mathcal{H})\to\Omega$ assigns an outcome sequence only for this context. The assignment is explicitly contextual/history-dependent and makes no claims about outcomes for other, incompatible PVMs. Hence $D$ is not a global valuation on the projection lattice, and the hypotheses of KS are not met.
@@ 4 Reversibility
Let us finally address briefly a natural question, that is: how large is, typically, the subset $S$ showing strong chain-transitivity? We provide a very simple result.
Theorem 4.7.
If $T:\mathbb{P}(\mathcal{H})\to\mathbb{P}(\mathcal{H})$ has no periodic points, then the system $(\mathbb{P}(\mathcal{H}),T)$ has a strongly chain-transitive, closed, invariant set $S\subseteq\mathbb{P}(\mathcal{H})$ of uncountable cardinality.
Proof.
By Theorem 4.5, a sufficient condition for having an infinite chain-transitive subsystem is that $T:\mathbb{P}(\mathcal{H})\to\mathbb{P}(\mathcal{H})$ has no periodic orbits, as no finite set of states can be chain-transitive unless there are periodic cycles.
Observe now that an infinite strong chain-transitive subset $S$ is necessarily uncountable. Indeed, assume by contradiction that $S$ is countable. Since $S$ is closed, it is a compact metric space. Any countable compact metric space contains an isolated point $v]\in S$. By invariance, $T([v])\in S$, and since $[v]$ is isolated, chain-transitivity implies that $[v]$ is periodic, contradicting the absence of periodic orbits. Hence $|S|>\aleph_{0}$. ∎
This purely topological argument is, of course, far from optimal. The existence of periodic orbits does not prevent at all the existence of an infinite chain-transitive subset, and the rough estimate by cardinality can be improved. The question of how large is the largest chain-transitive subsystem is explored in the forthcoming work [FLS_].
5 Conclusions
@@
Our work establishes a structural no-go for operational irreversibility along individual realized branches, under finite-dimensional, no-erasure conditions. Specifically, Theorems 4.5 and 4.6 tell us the following fact (which was indicated in the Introduction as our main result): in every finite-dimensional, open quantum system, there is a topologically closed, invariant (under the realization-induced dynamics) subsystem $S$ where any two states are connected via pseudo-orbits with arbitrarily small Fubini-Study error and energetic cost. By Def. 1.1, this means that $S$ is operationally reversible: no operational arrow of time arises.
@@
It is perhaps useful to frame this result by analogy with the logical structure of some classical thermodynamics concepts. The second law provides a universal lower bound on irreversibility, while Carnot’s theorem is an existence result showing that this bound can in fact be saturated: it identifies ideal cycles that achieve reversibility under suitable structural constraints. Landauer’s principle plays a similar role to the second law, including information in the entropy balance, furnishing a universal lower bound on dissipation for logically irreversible operations. Our result stands, structurally, in a similar relation to Landauer’s principle as Carnot’s theorem does to the second law: it is an existence statement identifying conditions (finite dimensionality and compactness) under which the bound is asymptotically saturated. Observe, however, that Carnot’s theorem is an inherently stronger result, within classical thermodynamics, than ours within quantum reversibility theory. Indeed, Carnot’s theorem provides a universal criterion, whereas our result only identifies a suitable subsystem showing quasi-reversibility.
@@
Let us finally remark that our argument crucially relies on the retention of the infinite record of outcomes provided by the realization map. This infinite history is essential because, to steer the system toward a target state with arbitrarily small FS error $\epsilon>0$, the perturbations must be tailored to the specific sequence of realized collapse outcomes encountered along the chain. Although only finitely many corrections are required to achieve any fixed $\epsilon$-precision, the number and nature of these corrections generally become unbounded as $\epsilon\to 0$, which means that, in principle, one needs the access to the entire future itinerary to prescribe the appropriate steering strategy at every scale. A similar observation is key in Bennett’s resolution of Maxwell’s demon paradox, where he showed that a demon’s finite memory forces logically irreversible erasure to accommodate new data, thereby incurring an entropic cost that preserves the second law [bennett1982thermodynamics]. Our no-erasure framework mirrors the idealized demon scenario, where an infinite memory sidesteps dissipation. Yet, as Bennett argued for any physical demon, real-world implementations face finite-size constraints that inevitably lead to erasure. Analogously, any practical realization of our procedure with a truncated record would introduce operational irreversibility, which recovers the second law.
@@
is a bijection between $\lambda$ and the set $\{x^{\gamma}\}_{0\leq\gamma<\lambda}$, so we have that $|\{x^{\gamma}\}_{0\leq\gamma<\lambda}|=|\lambda|$. By Hartogs’ Lemma [hartogs1915problem], we can take $\lambda$ so large that $|\{x^{\gamma}\}_{0\leq\gamma<\lambda}|>|X|$, which is a contradiction.